# The impact of gravitational waveform model systematics on the measurement of the Hubble Constant

### v1

@@ Abstract
Matching gravitational-wave observations of binary neutron stars with theoretical model predictions reveals important information about the sources, such as the masses and the distance to the stars. The latter can be used to determine the Hubble Constant, the rate at which the Universe expands. One general problem of all astrophysical measurements is that theoretical models only approximate the real underlying physics, which can lead to systematic uncertainties introducing biases. However, the extent of this bias for the distance measurement due to uncertainties of gravitational waveform models is unknown. In this study, we analyze a synthetic population of 38 binary neutron star sources measured with Advanced LIGO and Advanced Virgo at design sensitivity. We employ a set of four different waveform models and estimate model-dependent systematic biases on the extraction of the Hubble Constant using the bright siren method. Our results indicate that systematic biases are below statistical uncertainties for the current generation of gravitational wave detectors.
@@ I Introduction
Using gravitational-wave (GW) observations as a probe to infer the cosmic expansion rate was first proposed by Schutz in 1986 Schutz (1986). The GW emission of compact binary coalescences involving black holes or neutron stars can serve as cosmic distance indicators or so-called standard sirens and provide a measurement of the luminosity distance of the source. The landmark detection of GW170817 Abbott et al. (2017a, 2019), a binary neutron star (BNS) merger, by the LIGO and Virgo detectors Aasi et al. (2015a); Acernese et al. (2015a) along with its associated electromagnetic (EM) counterparts, GRB170817A and AT2017gfo (Abbott et al. (2017b) and refs. therein), marked a pivotal milestone in the field of cosmology. This event enabled the first direct measurement of the Hubble constant using a GW standard siren Abbott et al. (2017c). Although current constraints from GW observations are not yet able to resolve the Hubble tension, future observations could place further constraints on this cosmological parameter Chen et al. (2018); Feeney et al. (2019); Farr et al. (2019a); Feeney et al. (2021). Especially multi-messenger observations hold significant promise for improving the accuracy of Hubble constant determinations. This advantage arises from being able to combine the GW measurement of the distance to the source with an electromagnetic measurement of its redshift Abbott et al. (2017c); Guidorzi et al. (2017); Hotokezaka et al. (2019); Wang and Giannios (2021); Dietrich et al. (2020); Bulla et al. (2022); Palmese et al. (2024).
@@
While the standard siren methodology offers the advantage of not requiring re-calibration, as is common practice in the distance ladder method, there are different sources of systematic uncertainties that must be addressed. Previous studies such as Ref. Huang et al. (2022) investigated the impact of instrumental calibration uncertainties on Hubble constant measurements. They found that for single BNS events with a signal-to-noise ratio (SNR) of 50 for a network consisting of Advanced LIGO and Advanced Virgo operating at design sensitivity, the introduced bias is smaller than statistical uncertainties, but could accumulate as more events will be detected. Ref. Turski et al. (2023) examined the influence of galaxy redshift uncertainties using different galaxy redshift uncertainty models, with particular emphasis on photometric redshifts, which are known to exhibit substantial errors. Their findings indicate that while there may exist a potential bias, it remains smaller than statistical uncertainties. Furthermore, systematic uncertainties arising from the viewing angle of BNS mergers and EM selection effects can be major challenges Chen (2020); Chen et al. (2024); Mancarella et al. (2024). Building on Dhani et al. (2024), this work explores the systematic biases introduced by binary black hole (BBH) waveform models. The analysis reveals that these biases become more pronounced with increasing detector-frame total mass, binary asymmetry, and spin-precession effects. Notably, the study demonstrates on the basis of three high SNR events (golden events) that current waveform models lead to biased measurement of the Hubble constant, even for current detectors.
@@
In contrast to Ref. Dhani et al. (2024), our study delves into the issue of systematic uncertainties originating from BNS waveform models potentially influencing the determination of the distance to the source and, hence, the Hubble constant. Until now, this question remains, up to the best of our knowledge, unanswered. We analyze a synthetic population of 38 BNS sources, previously published in Ref. Kunert et al. (2022), with a set of four different waveform models and study model-dependent systematic errors. We employ the most recent Planck CMB measurement Aghanim et al. (2020) as a benchmark and investigate whether estimating $H_{0}$ using different models than the one used to generate the data introduces a systematic bias. As our results will demonstrate, GW model dependent biases have a minimal influence on the inferred value of the Hubble constant, for current detector networks.
@@
This paper is organized as follows. Section II provides details on modeling GW standard sirens, the simulated BNS source population, and on the Bayesian framework employed to extract estimates of $H_{0}$. Our results of combined $H_{0}$-estimates are presented in section III and section IV summarizes the key takeaways.
@@ II Methodology
II.1 Modelling standard sirens
@@ II.2 Population of synthetic BNS sources
In this study, we use the large set of mock data obtained in Ref. Kunert et al. (2022) and exploit the inferred posterior on the luminosity distance to study the impact of model uncertainties on the estimation of the Hubble Constant. Below, we recap the key elements of the injection study.
The population of $38$ synthetic binary neutron star sources were injected with stationary Gaussian noise into a network of interferometers assuming Advanced LIGO Aasi et al. (2015b) and Advanced Virgo Acernese et al. (2015b) design sensitivity. In light of current observations suggesting a uniform distribution for neutron star masses in gravitational-wave binaries Landry and Read (2021), we adopt a uniform prior sampled within $m_{1,2}\in[1,2]~{}M_{\odot}$. All sources were injected uniformly in co-moving volume with an optimal network SNR, $\rho$, ranging within $\rho\in[7,100]$ and lie within a maximum or threshold luminosity distance of $D^{\rm thr}_{L}=100$ Mpc. Hence, the population lies within a volume in which cosmology can be well modelled under the linear Hubble relation (for redshifts $z\ll 1$) given as
@@
in which $c$ is the speed of light. For all BNS signals in our population, we employed IMRPhenomD_NRTidalv2 (PhenDNRTv2) Dietrich et al. (2019a) as our injection model. We performed parameter estimation runs using parallel bilby Smith et al. (2020) for all sources in the population with four different models, which were TaylorF2 (TF2) Sathyaprakash and Dhurandhar (1991); Blanchet et al. (1995); Damour et al. (2001); Blanchet et al. (2004, 2005); Mishra et al. (2016); Mikoczi et al. (2005); Arun et al. (2009); Bohé et al. (2015); Mishra et al. (2016); Damour and Nagar (2010); Vines et al. (2011); Bini et al. (2012); Damour et al. (2012), IMRPhenomD_NRTidal (PhenDNRT) Husa et al. (2016); Khan et al. (2016); Taracchini et al. (2014); Dietrich et al. (2017, 2019b); Nagar et al. (2018); Bernuzzi et al. (2015), IMRPhenomD_NRTidalv2 (PhenDNRTv2), and SEOBNRv4_ROM_NRTidalv2 (SEOBNRTv2) Bohé et al. (2017a, b); Dietrich et al. (2019a) resulting in a total of 152 simulations. For further details on our injection setup, we refer the interested reader to Ref. Kunert et al. (2022).
@@ II.3 Bayesian framework
where $p(D_{L})$ is the prior on the luminosity distance and $p(\vec{\lambda})$ relates to priors for other GW parameters which are denoted as $\vec{\lambda}$. Analogous to the approach used in Ref.Abbott et al. (2017c), $N_{s}(H_{0})$ incorporates the selection effect imposed by the finite sensitivity of GW detectors and is
@@
integrated over the prior probability distributions of the parameters $D_{L},\vec{\lambda}$ and over “detectable” GW data sets. This factor is constructed on the assumption that the choice of whether or not an event is included in the analysis is a property purely of the observed data, $x_{\rm GW}$. In this analysis (as in Abbott et al. (2017c)) we assume that selection effects are dominated by the GW data, i.e., there are no selection effects for the EM counterpart. Additionally, we assume that GW detection is based on the observed SNR of the source [^1] In practice, we impose selection on the true, rather than the observed, luminosity distance of the source. This is the approach used widely in the literature. Selection on source parameters rather than observed data should really be treated differently Gair et al. (2023), but in practice this approximation does not change the results.. At the distances to which BNS events can be observed with current GW detectors, such a selection becomes a threshold in the luminosity distance of the source, which is what the GW detectors measure. As a consequence, the selection function has no dependence on the unknown cosmological parameters and so can be set to unity.
Since our dataset lacks direct redshift measurements for each BNS merger event, we need to model this information. We consider all EM observations as securely detected and, hence, neglect a correction for EM selection effects. In the absence of actual electromagnetic observational data for redshifts, we simulate a redshift measurement $\hat{z}$ for the $i$-th BNS merger event using a Gaussian probability density function
@@
This allows us to associate each BNS source with a simulated measured redshift sampled from a Gaussian distribution centered on its true, injected redshift value. We use the same Gaussian likelihood in Eq. (6) to obtain our Hubble constant posteriors. We examine the dependence of our analysis on the choice of the fractional error term, $C$, which contributes to the overall redshift uncertainty, $\sigma_{z}$, defined by the linear relationship $\sigma_{z}=z\cdot C$. This aims to quantify the sensitivity of our results to variations in the assumed fractional error value and, hence, in the precision of the EM measurement of the redshift. To infer joint constraints on $H_{0}$, we convert the measured redshift values to luminosity distances using Eq. (4).
@@ III Results
In accordance with the methodology detailed in section II.3, we obtained combined $H_{0}$-estimates for each GW model. As described, we aim to investigate the sensitivity of our results to variations in the fractional redshift error value $C$, which we explore for $C=0.5\%$ and $3\%$.
@@
First of all, we study how our estimate on $H_{0}$ changes with the number of successively combined GW signals for each BNS event for $C=0.03$, i.e., a 3% fractional error in the modeling of the observed redshift. The result is shown in Fig. 1 and shows that all models recover the Planck injection at $H_{0}=67.4~{}\rm kms^{-1}Mpc^{-1}$ within their uncertainty bands at 95% credible interval. Consistent with theoretical expectations, incorporating more GW events gradually decreases the uncertainties associated with the $H_{0}$ estimates. Among 38 combined GW events, PhenDNRTv2 (the “true” model) and PhenDNRT align best with the injected value, while SEOBNRTv2 and TF2 show slightly larger deviations from the injection. From Fig. 1, we find that systematic biases in the extraction of the Hubble constant are below statistical uncertainties.
@@
Figure 1: $H_{0}$-results versus the number of successively combined GW events for the GW models PhenDNRT (yellow), TF2 (purple), SEOBNRTv2 (orange), and the injection model PhenDNRTv2 (blue). Constraints on $H_{0}$ are reported at 95% credible interval. The injection value of the Planck measurement Aghanim et al. (2020) is shown as black dashed line. For this result, we modeled a $3$% uncertainty on the true, injected redshift as described in Sec. II.3.
@@
Secondly, we show the results for a 0.5% fractional error in Fig. 2, i.e., assuming more precise EM measurements of the observed redshift. We find that our injection model, PhenDNRTv2, recovers the injection value of the Planck measurement within 95% credible interval. Likewise, analyses employing the GW models PhenDNRT and TF2 recover the injection signal when considering the combined data from all BNS events. The only exception is the model SEOBNRTv2 which does not recover the injection when data from roughly 30 BNS events have been combined. Moreover, we find that there is a slight overestimation for all models when comparing with the injected Planck measurement, most likely due to statistical fluctuations.
@@
At this point, it is crucial to acknowledge that our recovered $H_{0}$ values presented in Fig. 2 depend upon the specific random realization of observed redshifts generated through Eq. (8). This implies that any variation in the random realization of redshift measurements for each event would lead to a corresponding change in the inferred $H_{0}$ posterior. While we present a single realization in Fig. 2, we investigated the dependence of our results with regard to different random realizations of the redshift measurement. Notably, while maintaining a fixed fractional error of 0.5%, we observed that the final, combined $H_{0}$ estimates exhibit a tendency to converge closer to the injected value under certain realizations (not shown in Fig. 2).
@@
In order to minimize the impact of stochastic redshift realizations and of specific BNS event orderings in our sample, we implement a double randomization procedure [^2] Despite this procedure, our GW posteriors are not randomized, which can lead to slight offsets or biases as we can see in our study.. This involves shuffling the BNS event order for 100 times and drawing 100 random redshift realizations. The result is shown in Fig. 3, for which we modeled a 0.5% uncertainty on the true, injected redshift. Our result depicts a convergence towards the injected Planck value for all GW models. Consistent with prior findings, it also reveals a progressive reduction in uncertainties and that all GW models recover the injection.
@@
Figure 2: $H_{0}$-results versus the number of successively combined GW events for the GW models PhenDNRT (yellow), TF2 (purple), SEOBNRTv2 (orange), and the injection model PhenDNRTv2 (blue). Constraints on $H_{0}$ are reported at 95% credible interval. The injection value of the Planck measurement Aghanim et al. (2020) is shown as black dashed line. For this result, we modeled a $0.5$% uncertainty on the true, injected redshift as described in Sec. II.3.
@@
From Figs. 2 and 3, we find that PhenDNRT tends to yield smaller $H_{0}$-values as compared to PhenDNRTv2. This could result from the fact that the PhenDNRTv2 includes an amplitude tidal correction term Dietrich et al. (2019a), while this is not the case for PhenDNRT. Due to this correction, the amplitude for PhenDNRT waveforms will be larger as compared to PhenDNRTv2. In order to compensate, PhenDNRT needs to go to higher luminosity distances and, hence, will predict smaller values for the Hubble constant. Another distinctive feature is that the uncertainty is smallest for TF2 when all data has been taken into account. To investigate this observation, we computed the Fisher information matrix (FIM) for both TF2 and PhenDNRTv2 across all events using GWfast Iacovelli et al. (2022a, b). The FIM analysis revealed that 55% of the events exhibited a lower distance error for TF2 as compared to PhenDNRTv2. This finding suggests that TF2 can achieve a narrower posterior width or lower variance. However, we would still expect it to have a larger bias, which would eventually dominate the uncertainty once sufficiently many events have been combined. Moreover, we repeat the analysis employing the Hubble constant measurement reported by the SHOES collaboration Riess et al. (2019) as the injected value, $H_{0}=74.03\pm 1.42\rm\ km\ s^{-1}Mpc^{-1}$. Figure 3 shows that we obtain consistent results compared to the scenario where we utilize the Planck measurement. Consequently, our primary conclusions remain unchanged.
@@
Overall, our results indicate that waveform model-dependent biases on the extraction of the Hubble constant are small for current detectors. Ascribing the subtle discrepancies observed in the combined $H_{0}$-estimates solely to specific model descriptions or assumptions presents a significant challenge. Firstly, GW models only possess a limited capability to accurately infer the distance and consequently, the Hubble constant, so we are dominated by statistical rather than systematic uncertainties. Secondly, our employed GW models exhibit interdependencies within specific regimes which hampers an unambiguous explanation attributable to individual models.
@@
Figure 3: $H_{0}$-results versus the number of successively combined GW events for the GW models PhenDNRT (yellow), TF2 (purple), SEOBNRTv2 (orange), and the injection model PhenDNRTv2 (blue) for both, using the SHOES Riess et al. (2019) measurement of $H_{0}=74.03\pm 1.42\rm\ km\ s^{-1}Mpc^{-1}$ and the Planck measurement Aghanim et al. (2020) of $H_{0}=67.4\pm 0.5\rm\ km\ s^{-1}Mpc^{-1}$ as injected value. We employed the same credible interval and fractional error on the redshift uncertainty as in Fig. 2, but randomly shuffled the order of BNS events in our sample for 100 times. We employ 100 random realizations of redshift measurements. The injection value of the Planck measurement is shown as black dashed line, while the injection of the SHOES measurement Riess et al. (2019) is shown as black dotted line.

### v2

# Impact of gravitational waveform model systematics on the measurement of the Hubble constant

@@ Abstract
Matching gravitational-wave observations of binary neutron stars with theoretical model predictions reveals important information about the sources, such as the masses and the distance to the stars. The latter can be used to determine the Hubble constant, the rate at which the Universe expands. One general problem of all astrophysical measurements is that theoretical models only approximate the real underlying physics, which can lead to systematic uncertainties introducing biases. However, the extent of this bias for the distance measurement due to uncertainties of gravitational waveform models is unknown. In this study, we analyze a synthetic population of 38 binary neutron star sources measured with Advanced LIGO and Advanced Virgo at design sensitivity. We employ a set of four different waveform models and estimate model-dependent systematic biases on the extraction of the Hubble constant using the bright siren method. Our results indicate that systematic biases are below statistical uncertainties for the current generation of gravitational-wave detectors.
@@ I Introduction
Using gravitational-wave (GW) observations as a probe to infer the cosmic expansion rate was first proposed by Schutz in 1986 Schutz (1986). The GW emission of compact binary coalescences involving black holes or neutron stars can serve as cosmic distance indicators or so-called standard sirens and provide a measurement of the luminosity distance of the source. The landmark detection of GW170817 Abbott et al. (2017a, 2019), a binary neutron star (BNS) merger, by the LIGO and Virgo detectors Aasi et al. (2015a); Acernese et al. (2015a) along with its associated electromagnetic (EM) counterparts, GRB170817A and AT2017gfo (Abbott et al. (2017b) and refs. therein), marked a pivotal milestone in the field of cosmology. This event enabled the first direct measurement of the Hubble constant using a GW standard siren Abbott et al. (2017c, 2021). Although current constraints from GW observations are not yet able to resolve the Hubble tension, future observations could place further constraints on this cosmological parameter Chen et al. (2018); Feeney et al. (2019); Farr et al. (2019a); Feeney et al. (2021). Multi-messenger observations especially hold significant promise for improving the accuracy of Hubble constant determinations. This advantage arises from being able to combine the GW measurement of the distance to the source with an electromagnetic measurement of its redshift Abbott et al. (2017c); Guidorzi et al. (2017); Hotokezaka et al. (2019); Wang and Giannios (2021); Dietrich et al. (2020); Bulla et al. (2022); Palmese et al. (2024).
@@
While the standard siren methodology offers the advantage of not requiring recalibration, as is common practice in the distance ladder method, there are different sources of systematic uncertainties that must be addressed. Previous studies such as Ref. Huang et al. (2022) investigated the impact of instrumental calibration uncertainties on Hubble constant measurements. They found that for single BNS events with a signal-to-noise ratio (SNR) of 50 for a network consisting of Advanced LIGO and Advanced Virgo operating at design sensitivity, the introduced bias is smaller than statistical uncertainties, but could accumulate as more events are detected. Reference Turski et al. (2023) examined the influence of galaxy redshift uncertainties using different galaxy redshift uncertainty models, with particular emphasis on photometric redshifts, which are known to exhibit substantial errors. Their findings indicate that while there may exist a potential bias, it remains smaller than statistical uncertainties. Furthermore, systematic uncertainties arising from the viewing angle of BNS mergers and EM selection effects can be major challenges Chen (2020); Chen et al. (2024); Mancarella et al. (2024). Building on Dhani et al. (2024), this work explores the systematic biases introduced by binary black hole (BBH) waveform models. The analysis reveals that these biases become more pronounced with increasing detector-frame total mass, binary asymmetry, and spin-precession effects. Notably, the study demonstrates on the basis of three high SNR events (golden events) that current waveform models lead to biased measurement of the Hubble constant, even for current detectors.
@@
In contrast to Ref. Dhani et al. (2024), our study delves into the issue of systematic uncertainties originating from BNS waveform models potentially influencing the determination of the distance to the source and, hence, the Hubble constant. Until now, this question remains, up to the best of our knowledge, unanswered. We analyze a synthetic population of 38 BNS sources, previously published in Ref. Kunert et al. (2022), with a set of four different waveform models and study model-dependent systematic errors. We employ the most recent Planck CMB measurement Aghanim et al. (2020) as a benchmark and investigate whether estimating $H_{0}$ using different models than the one used to generate the data introduces a systematic bias. As our results will demonstrate, GW model-dependent biases have a minimal influence on the inferred value of the Hubble constant for current detector networks.
@@
This paper is organized as follows. Section II provides details on modeling GW standard sirens, the simulated BNS source population, and on the Bayesian framework employed to extract estimates of $H_{0}$. Our results of combined $H_{0}$ estimates are presented in Sec. III and Sec. IV summarizes the key takeaways.
@@ II Methodology
II.1 Modeling standard sirens
@@ II.2 Population of synthetic BNS sources
In this study, we use the large set of mock data obtained in Ref. Kunert et al. (2022) and exploit the inferred posterior on the luminosity distance to study the impact of model uncertainties on the estimation of the Hubble constant. Below, we recap the key elements of the study.
The population of $38$ synthetic binary neutron star sources was simulated with stationary Gaussian noise into a network of interferometers assuming Advanced LIGO Aasi et al. (2015b) and Advanced Virgo Acernese et al. (2015b) design sensitivity. In light of current observations suggesting a uniform distribution for neutron star masses in gravitational-wave binaries Landry and Read (2021), we adopt a uniform prior sampled within $m_{1,2}\in[1,2]~{}M_{\odot}$. All sources were uniformly sampled in comoving volume with an optimal network SNR, $\rho$, ranging within $\rho\in[7,100]$ and lie within a maximum or threshold luminosity distance of $D^{\rm thr}_{L}=100$ Mpc. Hence, the population lies within a volume in which cosmology can be well modeled under the linear Hubble relation (for redshifts $z\ll 1$) given as
@@
in which $c$ is the speed of light. For all BNS signals in our population, we employed IMRPhenomD_NRTidalv2 (PhenDNRTv2) Dietrich et al. (2019a) as our reference model that we use for the creation of the simulated data. We performed parameter estimation runs using parallel bilby Smith et al. (2020) for all sources in the population with four different models, which were TaylorF2 (TF2) Sathyaprakash and Dhurandhar (1991); Blanchet et al. (1995); Damour et al. (2001); Blanchet et al. (2004, 2005); Mishra et al. (2016); Mikoczi et al. (2005); Arun et al. (2009); Bohé et al. (2015); Mishra et al. (2016); Damour and Nagar (2010); Vines et al. (2011); Bini et al. (2012); Damour et al. (2012), IMRPhenomD_NRTidal (PhenDNRT) Husa et al. (2016); Khan et al. (2016); Taracchini et al. (2014); Dietrich et al. (2017, 2019b); Nagar et al. (2018); Bernuzzi et al. (2015), IMRPhenomD_NRTidalv2 (PhenDNRTv2), and SEOBNRv4_ROM_NRTidalv2 (SEOBNRTv2) Bohé et al. (2017a, b); Dietrich et al. (2019a) resulting in a total of 152 simulations. For further details on our simulation setup, we refer the interested reader to Ref. Kunert et al. (2022).
@@ II.3 Bayesian framework
where $p(D_{L})$ is the prior on the luminosity distance, and other waveform parameters are denoted by $\vec{\lambda}$, similar to the notation employed in Ref. Abbott et al. (2017c). Hence, $p(\vec{\lambda})$ relates to priors for other GW parameters. Analogous to the approach used in Ref. Abbott et al. (2017c), $N_{s}(H_{0})$ incorporates the selection effect imposed by the finite sensitivity of GW detectors and is
@@
integrated over the prior probability distributions of the parameters $D_{L},\vec{\lambda}$ and over “detectable” GW datasets. This factor is constructed on the assumption that the choice of whether or not an event is included in the analysis is a property purely of the observed data, $x_{\rm GW}$. In this analysis (as in Abbott et al. (2017c)) we assume that selection effects are dominated by the GW data; i.e., there are no selection effects for the EM counterpart. Additionally, we assume that GW detection is based on the observed SNR of the source. [^1] In practice, we impose selection on the true, rather than the observed, luminosity distance of the source. This is the approach used widely in the literature. Selection on source parameters rather than observed data should really be treated differently Gair et al. (2023), but in practice this approximation does not change the results. At the distances to which BNS events can be observed with current GW detectors, such a selection becomes a threshold in the luminosity distance of the source, which is what the GW detectors measure. As a consequence, the selection function has no dependence on the unknown cosmological parameters and so can be set to unity.
Since our dataset lacks direct redshift measurements for each BNS merger event, we need to model this information. We consider all EM observations as securely detected and, hence, neglect a correction for EM selection effects. In the absence of actual electromagnetic observational data for redshifts, we simulate a redshift measurement $\hat{z}$ for the $i$th BNS merger event using a Gaussian probability density function
@@
This allows us to associate each BNS source with a simulated measured redshift sampled from a Gaussian distribution centered on its true redshift value. We use the same Gaussian likelihood in Eq. (6) to obtain our Hubble constant posteriors. We examine the dependence of our analysis on the choice of the fractional error term, $C$, which contributes to the overall redshift uncertainty, $\sigma_{z}$, defined by the linear relationship $\sigma_{z}=z\cdot C$. This aims to quantify the sensitivity of our results to variations in the assumed fractional error value and, hence, in the precision of the EM measurement of the redshift. To infer joint constraints on $H_{0}$, we convert the measured redshift values to luminosity distances using Eq. (4). In our study, we do not explore different redshift uncertainty models as demonstrated in Turski et al. (2023). Their work showed that the impact of different uncertainty models is much smaller than the current statistical errors. Hence, we expect that different redshift uncertainty models will not alter our results, which we will present in the next section.
@@ III Results
In accordance with the methodology detailed in Sec. II.3, we obtained combined $H_{0}$ estimates for each GW model. As described, we aim to investigate the sensitivity of our results to variations in the fractional redshift error value $C$, which we explore for $C=0.5\%$ and $3\%$. Our choice of the fractional error is driven by the inherent differences in redshift measurement techniques. Spectroscopic redshifts offer increased precision compared to photometric redshifts. Typical uncertainties for spectroscopic redshifts can be as low 0.01% Steinhardt et al. (2020), while photometric redshift uncertainties can range up to 10% or more Turski et al. (2023). Hence, we intend to mimic a more precise spectroscopic and a less precise photometric measurement of the redshift.
@@
First of all, we study how our estimate on $H_{0}$ changes with the number of successively combined GW signals for each BNS event for $C=0.03$, i.e., a 3% fractional error in the modeling of the observed redshift. The result is shown in Fig. 1 and shows that all models recover the Planck measurement at $H_{0}=67.4~{}\rm kms^{-1}Mpc^{-1}$ within their uncertainty bands at 95% credible interval. Consistent with theoretical expectations, incorporating more GW events gradually decreases the uncertainties associated with the $H_{0}$ estimates. Among 38 combined GW events, PhenDNRTv2 (the “true” model) and PhenDNRT align best with the Planck value, while SEOBNRTv2 and TF2 show slightly larger deviations. From Fig. 1, we find that systematic biases in the extraction of the Hubble constant are below statistical uncertainties.
@@
Figure 1: $H_{0}$ results versus the number of successively combined GW events for the GW models PhenDNRT (yellow), TF2 (purple), SEOBNRTv2 (orange), and the reference model PhenDNRTv2 (blue). Constraints on $H_{0}$ are reported at 95% credible interval. The reference value of the Planck measurement Aghanim et al. (2020) is shown as a black dashed line. For this result, we modeled a $3$% uncertainty on the true redshift as described in Sec. II.3.
@@
Second, we show the results for a 0.5% fractional error in Fig. 2, i.e., assuming more precise EM measurements of the observed redshift. We find that our reference model, PhenDNRTv2, recovers the Planck measurement within 95% credible interval. Likewise, analyses employing the GW models, PhenDNRT and TF2, recover this value when considering the combined data from all BNS events. The only exception is the model SEOBNRTv2 which does not recover the reference value when data from roughly 30 BNS events have been combined. Moreover, we find that there is a slight overestimation for all models when comparing with the Planck measurement. This is most likely due to a combination of statistical fluctuations and GW model-dependent differences in estimating the luminosity distance (as seen for TF2 in Fig. 1). The initial selection of GW events within our sample introduces another source of statistical fluctuation. However, the present bias is not significant for our conclusions as the true value lies within the posterior uncertainty.
@@
At this point, it is crucial to acknowledge that our recovered $H_{0}$ values presented in Fig. 2 depend upon the specific random realization of observed redshifts generated through Eq. (8). This implies that any variation in the random realization of redshift measurements for each event would lead to a corresponding change in the inferred $H_{0}$ posterior. While we present a single realization in Fig. 2, we investigated the dependence of our results concerning different random realizations of the redshift measurement. Notably, while maintaining a fixed fractional error of 0.5%, we observed that the final, combined $H_{0}$ estimates exhibit a tendency to converge closer to the reference value of the Planck measurement under certain realizations (not shown in Fig. 2).
@@
In order to minimize the impact of stochastic redshift realizations and of specific BNS event orderings in our sample, we implement a double randomization procedure. [^2] Despite this procedure, our GW posteriors are not randomized, which can lead to slight offsets or biases as we can see in our study. In this procedure, we draw $100$ random redshift values assuming a Gaussian distribution with a $0.5\%$ uncertainty for each BNS event within the sample. Equation (6) is then used to compute the combined $H_{0}$ value for $100$ times. In each iteration, the order of the BNS events is randomized, and a different drawn redshift value is used for each event. This procedure yields $100$ independent realizations of the combined $H_{0}$ posterior distribution. Finally, we compute the median of the distribution to extract the final results of $H_{0}$. The results are shown in Fig. 3, for which we modeled a $0.5\%$ uncertainty on the true redshift. Our result depicts a convergence towards the Planck value for all GW models. Consistent with prior findings, it also reveals a progressive uncertainty reduction and that all GW models recover the Planck measurement.
@@
Figure 2: $H_{0}$ results versus the number of successively combined GW events for the GW models PhenDNRT (yellow), TF2 (purple), SEOBNRTv2 (orange), and the reference model PhenDNRTv2 (blue). Constraints on $H_{0}$ are reported at 95% credible interval. The Planck measurement Aghanim et al. (2020) is shown as a black dashed line. For this result, we modeled a $0.5$% uncertainty on the true redshift as described in Sec. II.3.
@@
From Figs. 2 and 3, we find that PhenDNRT tends to yield smaller $H_{0}$ values as compared to PhenDNRTv2. This could result from the fact that the PhenDNRTv2 includes an amplitude tidal correction term Dietrich et al. (2019a), while this is not the case for PhenDNRT. Because of this correction, the amplitude for PhenDNRT waveforms will be larger as compared to PhenDNRTv2. In order to compensate, PhenDNRT needs to go to higher luminosity distances and, hence, will predict smaller values for the Hubble constant. Another distinctive feature is that the uncertainty is smallest for TF2 when all data has been taken into account. To investigate this observation, we computed the Fisher information matrix (FIM) for both TF2 and PhenDNRTv2 across all events using GWFAST Iacovelli et al. (2022a, b). The FIM analysis revealed that 55% of the events exhibited a lower distance error for TF2 as compared to PhenDNRTv2. This finding suggests that TF2 can achieve a narrower posterior width or lower variance. However, we would still expect it to have a larger bias, which would eventually dominate the uncertainty once sufficiently many events have been combined. Moreover, we repeat the analysis employing the Hubble constant measurement reported by the SHOES Collaboration Riess et al. (2019) as the reference value, $H_{0}=74.03\pm 1.42\rm\ km\ s^{-1}Mpc^{-1}$. Figure 3 shows that we obtain consistent results compared to the scenario where we utilize the Planck measurement. Consequently, our primary conclusions remain unchanged.
@@
Overall, our results indicate that waveform model-dependent biases on the extraction of the Hubble constant are small for current detectors. Ascribing the subtle discrepancies observed in the combined $H_{0}$ estimates solely to specific model descriptions or assumptions presents a significant challenge. First, GW models only possess a limited capability to accurately infer the distance and, consequently, the Hubble constant, so we are dominated by statistical rather than systematic uncertainties. Second, our employed GW models exhibit interdependencies within specific regimes, which hamper an unambiguous explanation attributable to individual models.
Finally, we use our results compared to the Planck measurement shown in Fig. 3 to estimate how many cumulative BNS events will be required to achieve a statistical uncertainty that falls below the waveform model systematics. To understand how much the final measurement of the Hubble constant might be affected by the choice of a certain model, we calculate the difference between the final $H_{0}$ value obtained using our reference model PhenDNRTv2 and the final values obtained using all the other models. The largest systematic shift of $\Delta H_{0}=0.18\rm\ km\ s^{-1}Mpc^{-1}$ is present when comparing to PhenDNRT, while the final result of TF2 has the smallest shift with $\Delta H_{0}=0.08\rm\ km\ s^{-1}Mpc^{-1}$. As the statistical uncertainty drops as $\sim 1/\sqrt{N_{\rm obs}}$, we can estimate the number of BNS events required to fall below the above-reported values. For our reference model PhenDNRTv2 to possess statistical uncertainties smaller than the systematic shift of $\Delta H_{0}=0.18(0.08)\rm\ km\ s^{-1}Mpc^{-1}$ roughly 3500 (16,500) BNS observations would be required with the current generation of GW detectors.
@@
Figure 3: $H_{0}$ results versus the number of successively combined GW events for the GW models PhenDNRT (yellow), TF2 (purple), SEOBNRTv2 (orange), and the reference model PhenDNRTv2 (blue) for both, using the SHOES measurement of $H_{0}=74.03\pm 1.42\rm\ km\ s^{-1}Mpc^{-1}$ (top panel)  Riess et al. (2019) and the Planck measurement of $H_{0}=67.4\pm 0.5\rm\ km\ s^{-1}Mpc^{-1}$ (bottom panel)  Aghanim et al. (2020) as reference values. We use the same credible interval and fractional error on the redshift uncertainty as in Fig. 2, but randomly shuffled the order of BNS events in our sample for 100 times, while employing 100 random realizations of redshift measurements.